answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are the possible outcomes if three coins are tossed and die is rolled?

There are 48 possible outcomes and I have no intention of listing them all.

They are all of the form CCCD

where C = H or T, and D takes the numeric values from 1 to 6.

If you draw 5 cards from a standard deck of cards what is the probability of drawing 5 hearts?

Probability (P)of drawing 5 hearts (H) is: P(H) and P(H) and P(H) and P(H) and P(H). There are 13 hearts in a 52 card deck. So, probability is: 13/52 * 12/51 * 11/50 * 10/49 * 9/48 or 154440/311875200 or 0.000495.

A couple has two children What is the probability that both are boys given that the first born is a boy?

The expected (but incorrect) answer to the question is 1/2.

It is not correct because:

  1. the probabilities of the gender of children in a family are not independent;
  2. the probability of a boy is not 1/2 but greater. For example, the global proportion of male births in 2012 was approx 52%.

What is the difference between standard deviation and mean?

The mean is the average value and the standard deviation is the variation from the mean value.

What is the difference between probability and membership function?

Membership is Not the same as probability.v Probability describes the uncertainty of event occurrence .The probability is an uncertainty associated with time.

vFuzziness describes event ambiguity , i.e., the degree to which an event occurs not whether it occurs

What is the probability that three coin flips will have at least one head?

Pr(3 flips at least one H) = 1 - Pr(3 flips, NO heads)

= 1 - Pr(3 flips, TTT) = 1 - (1/2)3 = 1 - 1/8 = 7/8

What are three famous events that occurred in 1969?

Concorde first flight

Robin Knox Johnson sails around the world, solo without stopping

The Apollo programme - first man on the moon

What is probability of getting head and tail alternatively in three throws of a coin?

If you mean 1: head 2:tail 3:head the probability is 1/2 x 1/2 x 1/2 = 1/8.

If you mean only that you want the heads and tails to alternate, that is, you mean the above or tail, head , tail, the probability is 1/8 + 1/8 = 1/4.

What is the probability of getting male and female offsring in human?

The global gender probability, at birth, is approx 0.517 for male and 0.483 for female. This is not the same as the gender ratio at conception because of gender-specific abortion.

How do you write 21 as a probability?

21 is a number. You cannot write that as a probability since it is not an event.

What is the probability of a black card?

50/50

or 1 in 2

as in could be either red or black

What is the probability of rolling an odd number less than 5?

If you're rolling an honest, normal cube, labeled 1 through 6 . . .

-- There are 6 possible outcomes of one roll.

-- Two of them (1 and 3) are odd numbers less than 5.

The probability is 2/6 = 1/3 = 331/3%.

The 'odds' are 2 to 1 against it.

What are some limitations of statistics?

limitations of statistics are as follows:

1. Statistics does not deal with an individual

2.It is not suitable to the study of qualitative phenomenon

3.Statistical relations are not exact

4.Statistics is liable to be misused

5.Statistics is only a means

How many people have the same birthday in the world?

It must be about 360,000 - the approximate number of world births per day
it is estematet to be that 150 million or billion(i cant remember) people share your birthday

Is grouped data better than ungrouped data?

Groued or ungrouped data all used to conclude something,but that conclusion is not 100% accurate,Whatever we use those just to live in this world,so dont bother about it . Groued or ungrouped data all used to conclude something,but that conclusion is not 100% accurate,Whatever we use those just to live in this world,so dont bother about it .

From the cross AA x AA the probability of producing a homozygous dominant offspring is?

From the cross Aa x Aa, the probability of producing a homozygous dominant offspring is 1/4 or (0.25).

How do you interpret Probability of default?

Distance-to-Default:-

- distance between the expected value of the asset and the default point

- after substitution into a normal c.d.f one gets probability of default

DD(t) =ln(V/F)+(µ-0.5*σ^2)*T/(σ*sqrt(T));

Where, V= value of the assets

F=Value of the liability/debt

µ= expected return of assets

σ=Volatility of the assets

T= Time

And Probability of default:-

PD(t) = NormDist(-DD)= Ɲ(-DD)

What two ways could a die be tested to see if its fair?

Two ways a die could be tested for fairness are empirical and analytic.

In the empirical method, you throw the die a large number of times and measure the distribution of the results - they should be evenly distributed within whatever level of tolerance (based on number of trials) that you want.

In the analytic method, you measure the dimensions of the die with extremely accurate equipment. This must include checks for physical symmetry and balance. Balance is checked with a balance beam that allows the die to be rotated about various axes to make sure the center of gravity is in the center of the die.

What is the probability of drawing 2 cards that are a 2 from a deck of 52?

The probability of drawing 2 cards that are two's from a standard deck of 52 playing cards is 1 in 221.

The probability of drawing the first two is 4 in 52 or 1 in 13. The probability of drawing the second two is 3 in 51. Multiply those two probabilities together and you get 3 in 663, or 1 in 221.

Trending Questions
Even numbered diamond cards are in a deck? If the letters of the word isosceles were put on cards and placed into a bag what is the probability that a vowel would be picked? Describe how to estimate the quotient of 4963 divided by 39? What is the possibility of something's being true or valid is known as? Relationship between relative frequency and probability of an event? How many different ways can you arrange 15 cars in 12 parking spaces? When a fair 6 sided dice is rolled what is the probability of scoring 2? How do you test theoretical probability? What is the first step in finding the probability of getting two numbers whose sum 2? What is the probability of getting a blackjack in blackjack? Can a goalie be replaced if he received a red card? What are the odds of hitting 2 straits in a row at 5 card stud? Does inferential statistics involve using a suitable population to draw a conclusion from a sample drawn from that population? What are the cards in Jake deck? How do you calculate retail when the cost and markup percent is given? What is the difference between independent events and mutually exclusive events? Did we at one time call flip flops go-heads? In a bag of 7 black marble 8 red marbles and 6 green marbles what is the least number of marbles you must pick without looking to guarantee two marbles of the same color? How do you calculate probability when mean and standards deviation are given? Which is better a score of 92 on a test with a mean of 71 and a standard deviation of 15 or a score of 688 on a test with a mean of 493 and a standard deviation of 150?